OpenStudy (anonymous):

witch is equivalent to 3 square root of 20 ^2 ???

OpenStudy (yanasidlinskiy):

\(\Huge\bf \color{yellow}{Welcome~to~OpenStudy!!}\hspace{-310pt}\color{cyan}{Welcome~to~OpenStudy!!}\hspace{-307.1pt}\color{midnightblue}{Welcome~to~\color{purple}{Open}}\color{blue}{Study!!!!}\)

OpenStudy (yanasidlinskiy):

Like this? \[3\sqrt{20^2}\]

OpenStudy (anonymous):

yes

OpenStudy (yanasidlinskiy):

Ok. have you had any similar problems to this before?

OpenStudy (anonymous):

no . ive nerver done one of these problems before idk where to start.

OpenStudy (yanasidlinskiy):

Ok. What is 20*20?

OpenStudy (anonymous):

400

OpenStudy (yanasidlinskiy):

Ok. So now Your problem looks like this: \[3\sqrt{400}\] What is the square root of 400? 20. Now you have: \[3\sqrt{20}\] opposite of dividing is multiplying. What does 20*3 =???

OpenStudy (anonymous):

60

OpenStudy (yanasidlinskiy):

Yep!:D Great Job!!!:D You got it!

OpenStudy (anonymous):

oh my god thank you so much !!

OpenStudy (yanasidlinskiy):

Your Welcome:D Don't forget to close the question and give a medal:D